# Beyond Image Prior: Embedding Noise Prior into Conditional Denoising Transformer

### v1

@@ 5 Conclusion and Discussions
In this paper, by rethinking the real image denoising task and revisiting the formation model of raw sensor noises, we have generalized a principle of the independence of image prior and noise prior, which guides an alternative conditional optimization to tackle the limitations of existing learning-based unconditional denoising methods. At the algorithmic level, we have presented a novel Condformer architecture, which effectively embeds the noise prior into the self-attention module. The noise prior is explicitly estimated using our LoNPE algorithm or network. Extensive experiments confirm the advantages of conditional optimization with noise prior, demonstrating that the proposed LoNPE and Condformer achieve superior performance on both synthetic and real noise statistics and image denoising tasks, respectively.
Nonetheless, due to the defective sensor and circuits technology, the formation model of raw sensor noise is actually more sophisticated than the Poisson-Gaussian noise model. For instance, read noise might follows a heavy-tailed Cauchy distribution [25] in extremely low-light environments, dark shading [26] can result from sensor non-uniformity, and more sophisticated noise models are emerging [85]. Therefore, future work should explore the estimation of noise prior in these more complex scenarios, aiming to further enhance noise estimation and denoising performance.

### v2

@@ 5 Conclusion and Discussions
In this paper, by rethinking the real image denoising task and revisiting the formation model of raw camera sensor noises, we have generalized a principle of the independence of image prior and noise prior. This principle guides an alternative conditional optimization to tackle the limitations of existing learning-based unconditional denoising methods. At the algorithmic level, we have presented a novel Condformer architecture, which effectively embeds the noise prior into the self-attention module. The noise prior is explicitly estimated using our LoNPE algorithm or network. Extensive experiments confirm the advantages of conditional optimization with noise prior, demonstrating that the proposed LoNPE and Condformer achieve superior performance on both synthetic and real noise statistics and image denoising tasks, respectively.
Nonetheless, there are other factors affecting imaging conditions that are relevant for noise statistics analysis, including aperture, sensor size/type and etc. Besides, due to the defective sensor and circuits technology, the formation model of raw sensor noise is actually more sophisticated than the Poisson-Gaussian noise model. For instance, read noise might follows a heavy-tailed Cauchy distribution 57 in extremely low-light environments, dark shading 17 can result from sensor non-uniformity, and more sophisticated noise models are emerging 8. Therefore, future work should explore the estimation of noise prior under different imaging factors and in more complex scenarios, aiming to further enhance noise estimation and denoising performance.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62202056, and the Fundamental Research Funds for the Central Universities under Grant 2243100002.
Data Availability
As introduced in Section 4.1.1, the datasets that support the findings of this study are openly available in the cited references. Our repository is provided in https://github.com/YuanfeiHuang/Condformer.